OpenStudy (anonymous):

Use graphical and numerical evidence to conjecture a value for the indicated limit. lim x -> infinity sign (x^2)/(2^x)

OpenStudy (anonymous):

@zepdrix can u help?

zepdrix (zepdrix):

Hmm what is the `sign` for?

OpenStudy (anonymous):

I read it as "infinity sign" -> \(\infty\)

OpenStudy (anonymous):

the infinity sing yes like space showed

OpenStudy (anonymous):

sign*

OpenStudy (anonymous):

\[\Large \lim_{x \to \infty} \frac{x^2}{2^x}\]

OpenStudy (anonymous):

so infinity sign with a negative on it yes space it that it

OpenStudy (anonymous):

yea space that the problem

zepdrix (zepdrix):

oh i see :) lol

OpenStudy (anonymous):

My english is a bit flawed, but does it include now a minus sign or not? I haven't investigated this limit yet, but I am sure it makes a big difference, because 2^x is a super fast growing function compared to x^2, hence it will reach infinity much faster then the numerator.

OpenStudy (anonymous):

no it doesn't include a minus sign

OpenStudy (anonymous):

\[\Large \lim_{x \to \infty} \frac{x^2}{e^{\ln2 x}} = \frac{\infty}{\infty} \] and then De L'hôpital. \[\Large \lim_{x \to \infty} \frac{2x}{\ln2e^{\ln 2 x}}= \lim_{x\to \infty} \frac{2}{(\ln2)^2e^{\ln2x}} =\frac{2}{\infty}=0\]

OpenStudy (anonymous):

so that how you conjecture the limit?

OpenStudy (anonymous):

Doubt that this counts as numerical evidence, it's rather analytical, but as for numeric, the runtime (time complexity) of 2^x is of a much higher order than 2^x, therefore it reaches infinity much faster then the numerator and that yields to zero.

OpenStudy (anonymous):

ok so how will i graph that information?

OpenStudy (anonymous):

graphical evidence means to draw or compute a graph, since they mention that option I guess you have a graphical calculator? If not, wolframalpha can do that for you.

OpenStudy (anonymous):

okay i'll try that. thanks

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

so look at my graph is it correct?

OpenStudy (anonymous):

@Spacelimbus

OpenStudy (anonymous):

ok i was wayy off.

OpenStudy (anonymous):

it's a complex graph.